Find a Matrix P that nationalizes A, and check your answer by computing PDP-1 1 0 0 0 1 -4 0 0 5
Answer #1

Consider, 0 stept DI-A= 16-18 o a-5 i characteristic ean is det CaI-A)=0 le. (ar)? (2-5) = 0 > a=1,1,5 are eigenvalues. Step-le. uzzo =) [e=05 & 44+42=0 Ty=-21 - E(s)= - Z z Honce, matrix P is 6 - - = -1 o o n D= o- = 1 ooh Clearly, PDP=A o where